The records of Hoffman Company reflected the following balances in the stockholders’ equity accounts at December 31, 2018:

Common stock, par $12 per share, 46,500 shares outstanding.
Preferred stock, 8 percent, par $12.5 per share, 7,210 shares outstanding.
Retained earnings, $233,000.


On January 1, 2019, the board of directors was considering the distribution of a $63,300 cash dividend. No dividends were paid during 2017 and 2018.


Required:

  1. Determine the total and per-share amounts that would be paid to the common stockholders and to the preferred stockholders under two independent assumptions:
  1. The preferred stock is noncumulative.
  2. The preferred stock is cumulative.
  1. Why were the dividends per share of common stock less for the cumulative preferred stock than the noncumulative preferred stock?
  2. What factors would cause a more favorable dividend for the common stockholders?

Determine the total and per-share amounts that would be paid to the common stockholders and to the preferred stockholders assuming the preferred stock is noncumulative. (Round "Per Share" to 2 decimal places and rest to the nearest dollar amount.)

Total Per Share
Paid to the Preferred Stockholders
Paid to the Common Stockholders

Determine the total and per-share amounts that would be paid to the common stockholders and to the preferred stockholders assuming the preferred stock is cumulative. (Do not round intermediate instructions. Round "Per Share" to 2 decimal places and rest to the nearest dollar amount.)

Total Per Share
Paid to the Preferred Stockholders
Paid to the Common Stockholders
Answer #1

Answer:

1a) Dividend distribution :
The preferred stock is non cumulative :
Total ($) Per Share ($)
Paid to preferred Stockholders 7210*12.5*8% =$ 7,210 7210/7210 = $1
Paid to Common Stockholders 63,300-7210 = $56,090 56,090/46,500 = $1.21
1 b) Dividend distribution :
The preferred stock is cumulative :
Total ($) Per Share ($)
Paid to preferred Stockholders 7210*12.5*8%*3Years = $21,630 21630/7210 = $3
Paid to Common Stockholders 63,300-21,630= $41,670 33813/46500=$ 0.90